# Leveraging Discrete Function Decomposability for Scientific Design

### v1

@@ Abstract
In the era of AI-driven science and engineering, we often want to design discrete objects in silico according to user-specified properties. For example, we may wish to design a protein to bind its target, arrange components within a circuit to minimize latency, or find materials with certain properties. Given a property predictive model, in silico design typically involves training a generative model over the design space (e.g., protein sequence space) to concentrate on designs with the desired properties. Distributional optimization—which can be formalized as an estimation of distribution algorithm or as reinforcement learning policy optimization—finds the generative model that maximizes an objective function in expectation. Optimizing a distribution over discrete-valued designs is in general challenging because of the combinatorial nature of the design space. However, many property predictors in scientific applications are decomposable in the sense that they can be factorized over design variables in a way that could in principle enable more effective optimization. For example, amino acids at a catalytic site of a protein may only loosely interact with amino acids of the rest of the protein to achieve maximal catalytic activity. Current distributional optimization algorithms are unable to make use of such decomposability structure. Herein, we propose and demonstrate use of a new distributional optimization algorithm, Decomposition-Aware Distributional Optimization (DADO), that can leverage any decomposability defined by a junction tree on the design variables, to make optimization more efficient. At its core, DADO employs a soft-factorized “search distribution”—a learned generative model—for efficient navigation of the search space, invoking graph message-passing to coordinate optimization across linked factors.
@@ 1 Design in discrete state spaces
Figure 1: Key components of DADO. a, DADO requires as input an objective function in its decomposed form, $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$, which corresponds to a junction tree. Here we show a junction tree with nodes of size 1, i.e., a regular tree, for simplicity. Variables with edges interact together to directly influence $f$. Some variables participate in multiple component functions, requiring coordination in the form of message-passing. b, To update the search distribution at each iteration, naive EDAs weight entire samples drawn from a joint distribution over all design variables by scoring with $f(x)$. Shade of green denotes more to less optimal scores. In contrast, DADO leverages the decomposition of $f(x)$ to weight samples in a more local manner, according to the decomposition. Specifically, DADO uses message-passing to compute value functions that account for $x_{i}$ interacting with its descendants. Correspondingly, the value functions serve as the weights for each part of the search distribution, which is factorized like $f$. Optional shaping function $W$ is omitted for clarity. c, Example performance comparison on a synthetic problem with an exact tree decomposition over a discrete design space of size $20^{50}$ ($D=20,L=50$). Each of the two methods drew 100 samples per iteration. We evaluated these samples with $f(x)$, computing the per-iteration mean and 95% confidence interval. Results shown were averaged over 20 random seeds for the same $f(x)$ (details in Sec. 4.1, Fig. 2). The p-value shown is from a two-sided paired t-test that AUC of the per-iteration means is different between methods, using the 20 mean curves.
@@
The integration of AI into scientific research has opened new avenues for property-driven, in silico design of discrete objects—from molecular structures like proteins, to engineered systems like circuits—where computational methods guide design with user-specified properties. For example, we may aim to design an amino acid sequence for a protein so that the protein binds its target.
@@
Given a property predictive model, $f(x)$, the simplest version of in silico design entails enumerating all possible designs, $x=[x_{1},x_{2},...,x_{L}]\in\mathcal{X}$ (e.g., all possible amino acids of length $L$), evaluating each one under the predictive model for property, $s=f(x)$, and choosing the design with the highest $s$. Such a setup is complicated by two primary challenges. First, in most realistic problems, the design space is too large to fully enumerate. Second, if $f(x)$ has parameters estimated from data, then it is most likely not accurate over the whole space. While much work has attempted to address the second problem (e.g.brookes19_cbas; trabucco21_coms; uehara24_braid), little attention has been placed on the first one in the context of decomposable predictive models. Herein, we focus on this issue: performing efficient optimization in high-dimensional discrete design spaces that can be decomposed according to structure in $f(x)$. As the second problem is complementary to ours, we do not here consider it—solutions for each can be implemented together. We focus our examples and experiments on designing amino acid sequences, but our proposed method is general, applicable to design on any discrete space. For a $D$-amino acid alphabet, where typically $D=20$, finding the highest-scoring design, $x^{*}$, naively requires $D^{L}$ evaluations of $f(x)$. To develop some intuition, in this paragraph only, assume that $D=2$ so that $x$ is binary. First, consider the simple but unrealistic case of full linear additivity, where $f(x)=\beta_{1}x_{1}+\beta_{2}x_{2},\ldots,\beta_{L}x_{L}$, for scalar parameters $\{\beta_{i}\}$. In such a scenario, one can find $x^{*}$ by optimizing each of the $L$ components of $x$ separately because they do not interact with each other at all, requiring evaluating only $D\times L\ll D^{L}$ partial designs. A more realistic setting would allow for more complicated functional forms while still having some notion of linear additivity. For example, consider the form $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$, where $C_{i}$ denotes an arbitrary “component” function on a set of design variables, $\hat{x}_{i}$ (e.g., $\hat{x}_{i}=\{x_{1},x_{3},x_{8}\}$), such as $C_{1}(\hat{x}_{1})=C_{1}(x_{1},x_{3},x_{8})=\exp(7x_{1}x_{3}-3x_{3}x_{8})$. As long as sets of variables don’t overlap, i.e., $x_{i}$ only appears in a single component function $C_{j}$, we can separately optimize each component $C_{j}$, which is over fewer variables than $f(x)$, resulting in an easier and more efficient optimization. In the most general setting, encompassing any function on discrete variables, sets of variables can overlap across component functions, thereby tying them together and making the optimization correspondingly harder—but as we will see, still much easier than naive optimization of $f$ in many cases (Fig. 1). This difference in optimization difficulty between the decomposed and naive scenarios increases with the alphabet size and number of design variables, which are both large in many realistic settings. If $f$ has no decomposability at all, it can be represented as a single component function over all variables, namely, itself, in which case there is no efficiency to be gained. However, in most real problems, we expect some level of decomposability because the world is generally structured. For example, in designing a protein to bind to a target, a few key sequence positions may make up the binding interface, which jointly, primarily dictate the binding strength, whereas the other positions may be involved, largely separately from the binding site, in stabilizing the protein for binding. Our approach will not require full decomposability so that these types of factors can still interact with each other. We refer to this as “soft-decomposability”, whereby some variables can participate in multiple component functions.
To optimize a soft-decomposable function, one can employ classical message-passing algorithms, which leverage decomposability to exactly find the global optimum with the lowest possible time complexity (vlassis04_maxplus). However, for reasons discussed momentarily, we are interested in distributional optimization (DO), wherein a standard optimization problem over the design space, $x^{*}=\arg\max_{x\in\mathcal{X}}f(x)$, is replaced by one over the parameters of a generative model, $p_{\theta}(x)$, namely, $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{p_{\theta}(x)}\left[f(x)\right]$. These two optimization problems are equivalent in that the optimum of each is the same when $p_{\theta}(x)$ has capacity to place all of its mass on the maximal designs. Instead of naive, decomposition-unaware DO, we will introduce a decomposition-aware version of DO that takes advantage of soft-decomposability in $f$ by way of a soft-factorized $p_{\theta}(x)$, each factor of which is responsible for a single component of $x$ (e.g., $\hat{x}_{i}$). This factorization results in an overall effective search space that is substantially smaller than $\mathcal{X}$, like in our examples above. Moreover, we will show that each factor distribution can be updated separately from all others, such that given a fixed sampling budget for training $p_{\theta}(x)$, each factor can apply the full budget to its smaller subset of design dimensions. As a result, our decomposition-aware DO algorithm will be more statistically efficient than naive DO and thus navigate toward promising regions of $\mathcal{X}$ more quickly.
@@
Our interest in the distributional optimization formulation is motivated by its extensibility. First, DO enables one to directly use innovations from the Estimation of Distribution Algorithm (EDA) (larranaga01_eda; brookes20_eda_em) and policy optimization (peters07_rwr; peng19_awr) communities. Of particular note are methods that enable combining a pre-trained, unconditional generative model, $p(x)$, with a property predictor, $p(y|x)$, to execute Bayes rule so as to obtain a sampling distribution that can be used for design, $p(x|y\in Y)$ (e.g.brookes19_cbas; Fannjiang2020; uehara24_braid). Second, as the key object that navigates the search space, $p_{\theta}(x)$, is a generative model, we stand to benefit from advances in generative modeling. Herein, for clarity of contribution, we focus on the purest form of the EDA, without entropy regularization, a prior, or conditioning, leaving such extensions to future work.
@@
We develop a DO algorithm in the form of a generalized EDA/policy optimization algorithm that can leverage any decomposability in $f(x)$ in the form described earlier as $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$ (also equivalently described by a junction tree), to more efficiently navigate the design space and find designs with high $f(x)$ (Fig. 1). We call our method Decomposition-Aware Distributional Optimization (DADO). We first empirically investigate DADO on entirely synthetic problems, illustrating that DADO indeed yields designs with high $f(x)$ more quickly than a version that is not decomposition-aware. Next we investigate DADO for problems anchored on real protein data, further substantiating that leveraging decomposition helps.
@@
As DADO requires as input a soft-decomposed objective, $f(x)$, for our protein problems we construct these as follows. We first obtain a graph topology of which variables (protein positions) are connected by thresholding residue distances from an AlphaFold3 structure, and adding an edge between pairs of positions within this threshold (abramson24_af3; brookes22_sparsity). From this graph we can construct a junction tree, which in turn directly maps to the functional form of the predictive model. Recall that junction trees represent arbitrarily complex relationships between variables by transforming any graph on those variables into a tree structure, where each node contains a set of the original variables, and where these sets may be overlapping in adjacent nodes (lauritzen88_JT). Given the junction tree topology, we can then fit a neural network adherent to the topology on sequence-property pairs, $(x^{k},y^{k})$. More precisely, each component $C_{i}(\hat{x}_{i})$ is parameterized and the sum of all components, $f(x^{k})=\sum_{i}^{\kappa}C_{i}(\hat{x}^{k}_{i})$, regressed onto property labels $y^{k}$. Analogous procedures in different application areas could include the following, for example. In circuit design, certain topologies of components may be prescribed ahead of time according to production constraints or domain knowledge. When designing a telescope, designers consider different arrangements of lenses—i.e., topologies—and optimize the diameter, curvature, material, coating, etc. for each component.
2 Decomposition-Aware Distributional Optimization (DADO)
@@
We begin by reminding the reader of the three primary steps that are iterated in a naive EDA, which is unaware of any decomposability (e.g.brookes20_eda_em). From there, we describe how infusing this algorithm with awareness of the known decomposition of $f(x)$ stands to make the EDA more statistically efficient, after which we formally introduce DADO.
@@
Intuitively, one can think of an EDA as having a search distribution, $p_{\theta}(x)$, that acts as a spotlight on the design space, which is iteratively moved toward areas of the space with high $f(x)$ in expectation. Modern EDAs parameterize the search distribution with neural network generative models, such as a Variational Autoencoder (kingma13_vae; e.g.,  brookes19_cbas). In a decomposition-unaware EDA, after having initialized the search distribution, $p_{\theta^{0}}(x)$, with initial parameters, $\theta^{0}$, the EDA proceeds by iterating through these three steps for $n=1\dots N$ or until some convergence criteria has been met ( larranaga01_eda; brookes20_eda_em; Sec. A.2):
@@
The weighted MLE problem in step 3 is typically solved with gradient descent; we will choose to use only a fixed number of gradient steps, which can be theoretically justified through the equivalence of EDAs to Expectation-Maximization (brookes20_eda_em).
Now that a regular EDA is fresh in our minds, we can consider what it might mean to infuse it with knowledge of how $f(x)$ is decomposed, and how this might prove useful. Before discussing the general case, here we first build intuition by considering a simpler example, one that is “fully”-decomposable rather than only soft-decomposable. That is, consider $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2})$ where the two meta-variables $\hat{x}_{1}$ and $\hat{x}_{2}$ do not share any original variables, $x_{i}$, corresponding to a junction tree without any edges, only independent nodes. In this setting, we can break the EDA down into two independent EDAs, each with reduced dimensionality compared to the original EDA. Specifically, we can employ a fully-factorized search distribution (“fully” again denoting that no variable appears in more than one factor), $p_{\theta}(x)=p_{\theta_{1}}(\hat{x}_{1})p_{\theta_{2}}(\hat{x}_{2})$. We modify the three steps of the EDA as follows:
  1. 1.
  2. 2.
  3. 3.
Earlier we described decomposability as $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$ to help provide some intuition. However, to develop DADO it will be more convenient to equivalently describe the decomposability with a junction tree defined by nodes that contain multiple design variables and undirected edges that connect the nodes. We will define this junction tree decomposition in Sec. 2.1 and explain its connection to the above decomposability in Sec. A.1. It suffices to say here that design variables $x_{i}$ that occurred uniquely in one meta-variable $\hat{x}_{j}$ now sit in a single node, $\tilde{x}_{k}$; variables that previously occurred in multiple meta-variables now appear in multiple adjacent nodes. With the idea of a junction tree in hand, we are now set to complete our intuitive overview of DADO, which will use a factorized search distribution matching the topology of the junction tree. Namely, we will convert the undirected junction tree into a directed, acyclic graph (DAG) by rooting it and then directing the edges outward from the root. This operation will yield a search distribution that can be written as $p_{\theta}(x)=p_{\theta}(\tilde{x}_{r})\sum_{i}{p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})}$, for root node $\tilde{x}_{r}$ and non-root nodes, $\tilde{x}_{i}$, with parents, $\tilde{x}_{\text{p}(i)}$. Note that each non-root node has exactly one parent in a DAG. Each search distribution factor is a conditional probability distribution that will be parameterized by a neural network such as an autoregressive model.
  1. 1.
  2. 2.
  3. 3.
Having provided an intuitive overview of DADO (Fig. 1), we next provide a more formal exposition.

### v2

@@ Abstract
In the era of AI-driven science and engineering, we often want to design discrete objects (e.g., circuits, proteins, materials) in silico according to user-specified properties (e.g., that a protein binds its target). Given a property predictive model, in silico design typically involves training a generative model over the design space (e.g., over the set of all length-$L$ proteins) to concentrate on designs with the desired properties. Distributional optimization, formalized as an estimation of distribution algorithm or as reinforcement learning policy optimization, maximizes an objective function in expectation over samples. Optimizing a distribution over discrete-valued designs is in general challenging due to the combinatorial nature of the design space. However, many property predictors in scientific applications are decomposable in the sense that they can be factorized over design variables in a way that will prove useful. For example, the active site amino acids in a catalytic protein may need to only loosely interact with the rest of the protein for maximal catalytic activity. Current distributional optimization algorithms are unable to make use of such structure, which could dramatically improve the optimization. Herein, we propose and demonstrate use of a new distributional optimization algorithm, Decomposition-Aware Distributional Optimization (DADO), that can leverage any decomposability defined by a junction tree on the design variables. At its core, DADO employs a factorized “search distribution”—a learned generative model—for efficient navigation of the search space, and invokes graph message passing to coordinate optimization across all variables.
footnotetext: 1Department of Electrical Engineering and Computer Sciences, UC Berkeley. 2Center for Computational Biology, UC Berkeley. 3UC Berkeley–UCSF Graduate Program in Bioengineering, UC San Francisco.
@@ 1 Design in Discrete State Spaces
The integration of AI into scientific research has opened new avenues for property-driven, in silico design of discrete objects—from molecular structures like proteins, to fabricated systems like circuits—where computational methods guide design with user-specified properties. For example, we may seek to design an amino acid sequence for a protein so that the protein binds its target.
Given a property predictive model, $f(x)$, the simplest version of in silico design entails enumerating all possible designs, $x=[x_{1},x_{2},...,x_{L}]\in\mathcal{X}$ (e.g., all possible amino acid sequences of length $L$), evaluating each one under the predictive model, $s=f(x)$, and choosing the design with the highest $s$. Such a setup is complicated by two primary challenges. First, in most realistic problems, the design space is too large to fully enumerate. Second, if $f(x)$ has parameters estimated from data, then it is most likely not accurate over the whole space. While many works address the second problem, there has been little recent development on the first; as such, we focus herein on performing efficient optimization in high-dimensional discrete design spaces. The second problem can be handled with approaches complementary to ours (e.g.Brookes et al. (2019); Trabucco et al. (2021); Uehara et al. (2024)), and is not here considered. We will focus our examples and experiments on designing amino acid sequences, but our method is general—applicable to design on any discrete space.
@@
Finding the highest-scoring design, $x^{*}$, naively requires $D^{L}$ evaluations of $f(x)$ for a $D$-amino acid alphabet (typically $D=20$). To develop some intuition, in this section only, let us assume that $D=2$ so that $x$ is a binary sequence. First, consider the simple but unrealistic case where $f(x)=\beta_{1}x_{1}+\beta_{2}x_{2},\ldots,\beta_{L}x_{L}$, for scalar parameters $\{\beta_{i}\}$. In such a scenario, finding the design with the highest $f(x)$ requires considering only $D\times L$ partial designs because the $L$ components of $x$ do not interact with each other, acting only linearly additively. A more realistic setting would allow for more complicated functional forms while maintaining some notion of linear additivity. For example, consider the form $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$, where $C_{i}$ denotes an arbitrary function on a set of design variables, $\hat{x}_{i}$, such as $C_{1}(\hat{x}_{1})=C_{1}(x_{1},x_{3},x_{8})=\exp(7x_{1}x_{3}-3x_{3}x_{8})$. In the case where sets of variables do not overlap, i.e., where $x_{i}$ can appear in only one component $C_{j}$, finding the global optimum of $f(x)$ requires a number of evaluations of component functions that scales as $D^{M}$, where $M\leq L$ is the cardinality of the largest variable set. When the sets of variables do overlap across component functions, tying them together, then the number of required evaluations becomes correspondingly higher. Note that this more general form of linear additivity admits representation of any function, possibly requiring that $M=L$ in which case there is no linear additivity, nor consequently, decomposability. In realistic settings with many design variables and a larger alphabet (e.g., $D=20$), the difference in number of evaluations required between the decomposed and standard scenarios is even greater still. Importantly, in most real problems, we expect some level of decomposability. For example, in designing a protein to bind to a target, a few key sequence positions may make up the binding interface, which primarily dictates the binding strength, whereas the other positions may be involved in stabilizing the protein for binding.
Classical message passing algorithms can leverage the aforementioned general type of structure in $f(x)$ to exactly find the global optimizer with the lowest possible time complexity (Vlassis et al., 2004). However, for reasons discussed momentarily, we are interested in distributional optimization, wherein a standard optimization problem over the design space, $x^{*}=\arg\max_{x\in\mathcal{X}}f(x)$, is replaced by one over the parameters of a generative model, $p_{\theta}(x)$, namely, $\theta^{*}=\arg\max_{\theta}\mathbb{E}_{p_{\theta}(x)}\left[f(x)\right]$. These formulations are equivalent in that the optimum of each is the same, assuming that $p_{\theta}(x)$ has the capacity to place all of its mass on $x^{*}$. However, each formulation lends itself to different algorithms and extensions, mentioned below. Distributional optimization may employ strategies to prevent $p_{\theta}(x)$—the search distribution/policy—from collapsing to a point mass, such as by using a prior (Brookes et al., 2019) or entropy regularizer (Ziebart et al., 2008).
@@
Our interest in the distributional optimization formulation is motivated by its extensibility. First, such a setup enables us to directly use innovations from the Estimation of Distribution Algorithm (EDA) (Brookes et al., 2020; Larrañaga and Lozano, 2001) and policy optimization (Peters and Schaal, 2007; Peng et al., 2019) communities. Of particular note are methods that enable combining a pre-trained, unconditional generative model, $p(x)$, with a property predictor, $p(y|x)$, to execute Bayes rule and thereby obtain a sampling distribution that can be used for design, $p(x|y\in Y)$ (e.g., Brookes et al. (2019); Fannjiang and Listgarten (2020); Uehara et al. (2024)). Second, as the key object that navigates the search space, $p_{\theta}(x)$, is a generative model, we stand to benefit from advances in generative modeling. Herein, for clarity of contribution, we focus on the purest form of the EDA, without entropy regularization or a prior. We leave such extensions to future work.
@@
We develop a distributional optimization algorithm in the form of a generalized EDA/policy optimization algorithm that can leverage any decomposability in $f(x)$ in the form described earlier as $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$, to more efficiently navigate the design space and find high-performing designs quickly (Fig. 1). We call our method Decomposition-Aware Distributional Optimization (DADO). We first empirically investigate DADO on synthetic examples, illustrating that the anticipated optimization efficiency emerges compared to decomposition-unaware baselines. Then we further substantiate this efficiency on problems anchored on real protein data, that is, by optimizing protein property predictive models. Additionally, in case studies on a few protein predictive models, we find that accuracy is robust to modifications of the decomposition, suggesting that perfect a priori knowledge of decomposability is not required.
@@
As DADO requires a decomposed objective of the form $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$, on our protein data problems, we construct these as follows. We first obtain a graph topology of which variables (protein residues) are coupled by thresholding residue distances from an AlphaFold3 structure (Abramson et al., 2024; Brookes et al., 2022). This graph dictates the functional form of the predictive model, by way of an automatically constructed junction tree. Recall that junction trees represent arbitrarily complex relationships between random variables by transforming any graph into a tree structure, where each node contains a set of the original variables, and where these sets may be overlapping in adjacent nodes (Lauritzen and Spiegelhalter, 1988). Then we fit a neural network that enforces this decomposition, to the training data. Analogous procedures in different application areas could include the following, for example. In circuit design, certain topologies of components may be prescribed ahead of time according to production constraints or domain knowledge. When designing a telescope, designers consider different arrangements of lenses—i.e., topologies—and optimize the diameter, curvature, material, coating, etc. for each component.
Figure 1: Key components of DADO. a, DADO requires as input an objective function in its decomposed form, $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$, which corresponds to a junction tree. Here we show a junction tree with nodes of size 1, i.e., a regular tree, for simplicity. Variables with edges interact to directly influence $f$. Some variables participate in multiple component functions, requiring coordination. b, To update the search distribution at each iteration, naive EDAs weight entire samples, drawn from a joint distribution over all design variables, with $f(x)$. In contrast, DADO leverages the decomposition of $f(x)$ to separately weight sample dimensions for each node. DADO uses message-passing to compute value functions, $Q_{i}$, accounting for each $x_{i}$ interacting with its descendants. These $Q_{i}$ serve as weights for each part of the search distribution, which is factorized like $f$. c, Example performance comparison on a synthetic problem with an exact tree decomposition over a discrete design space of size $20^{50}$ ($D=20,L=50$). Each method drew 100 samples per iteration. We evaluated these samples with $f(x)$, computing the per-iteration mean and 95% confidence interval. Results shown were averaged over 20 random seeds for the same $f$ (details in Sec. 4). The p-value is from a two-sided paired t-test that the mean at the final iteration is different between methods, over the 20 seeds.
2 Decomposition-Aware Discrete Optimization (DADO)
@@
We begin by reminding the reader of the three primary steps that are iterated in a standard EDA, which is unaware of any decomposability (e.g., Brookes et al. (2020)). From there, we describe how infusing this algorithm with awareness of the known decomposition of $f(x)$ stands to make the EDA more statistically efficient, after which we formally introduce DADO.
@@
Intuitively, one can think of the search distribution as a spotlight on the design space, which iteratively gets moved toward areas with high $f(x)$ in expectation. Modern EDAs parameterize the search model with neural network generative models, such as a Variational Autoencoder (e.g.,  Brookes et al. (2019)). In a standard EDA, after initializing the search distribution, $p_{\theta^{0}}(x)$, the EDA proceeds by iterating through these three steps either $N$ times or until convergence (Sec. A.2):
@@
    Update search distribution parameters by weighted maximum likelihood estimation (MLE), using the weighted samples, $\theta^{n+1}=\arg\max_{\theta}\;\mathbb{E}_{\{x^{k}\}}w^{k}\log p_{\theta}(x^{k})$.
@@
The weighted MLE problem is typically solved with gradient descent; choosing to use only a fixed number of gradient steps, as we will do, can be theoretically justified through the equivalence of EDAs to Expectation-Maximization (Brookes et al., 2020).
Now that a standard EDA is fresh in our minds, we can consider what it might mean to infuse it with knowledge of how $f(x)$ is decomposed, and how this might prove useful. Before discussing the general case, let us first build intuition with a simpler example. Specifically, consider $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2})$ where there is no overlap in variables, $x_{i}$, between the meta-variables $\hat{x}_{1}$ and $\hat{x}_{2}$. In this setting, we can replace step 1, of sampling from one search model over all the design variables, $x=[\hat{x}_{1},\hat{x}_{2}]=[x_{1},x_{2},...,x_{L}]\sim p_{\theta^{n}}(x)$, with instead sampling each meta-variable from its own search model separately, $\hat{x}_{1}^{k}\sim p_{\theta^{n}_{1}}(\hat{x}_{1})$ and $\hat{x}_{2}^{k}\sim p_{\theta^{n}_{2}}(\hat{x}_{2})$. Having done so, we can execute steps 2 and 3 in a similar manner by scoring each meta-variable separately with its component function (e.g., $w_{1}^{k}=C_{1}(\hat{x}_{1}^{k})$) and then updating each search model independently of the other, via weighted MLE, to obtain $p_{\theta^{n+1}_{1}}(\hat{x}_{1})$ and $p_{\theta^{n+1}_{2}}(\hat{x}_{2})$ (see Sec. A.3 for side-by-side algorithm sketches).
What has this bought us? We have split one optimization problem over a combinatorial space into two independent optimization problems, each over a much smaller space. For a problem where each meta-variable comprises just 15 binary design variables, we have transformed an optimization over a space of size $2^{30}\approx 10^{9}$ to two problems each of size $2^{15}\approx 10^{5}$. For larger alphabets and sequence lengths, these differences will be larger still. Importantly, the weighted MLE update with a fixed number of samples, $K$, is more statistically efficient for each lower-dimensional search model than it is for a decomposition-unaware EDA using a single joint search model over all design variables.
The general case allows overlapping variables across component functions, requiring coordination between meta-variables—we cannot divide and conquer as above; rather, we must divide, collaborate, and then conquer. It will be necessary to equivalently describe the decomposability of $f(x)$ with a junction tree, defined by nodes corresponding to meta-variables and undirected edges connecting them (details in Sec. 2.1 and Sec. A.1). Given such a junction tree decomposition, we use a factorized search distribution—a directed, acyclic graphical (DAG) model—matching its topology. Each factor specifies the density of a node’s variables conditioned on its parent node’s variables, and will be parameterized by a neural network such as an autoregressive model. Sampling from such a distribution is both computationally efficient and easy, so step 1 of the EDA remains straightforward. However, steps 2 and 3 of the standard EDA are not so easily generalized. To update the factorized search distribution in a manner that reduces the effective size of the optimization problem, we will need to generalize “max-plus” message passing, which yields a global optimum $x^{*}=\arg\max_{x\in\mathcal{X}}f(x)$ (Vlassis et al., 2004), to a procedure integrating message passing with a factorized search model fitting step, thus constituting a decomposition-aware EDA. Having provided an intuitive overview of DADO (Fig. 1), we now proceed to a more formal exposition.